You would like to create a portfolio that is equally invested in a risk-free asset and two stocks. The one stock has a beta of .75. What does the beta of the second stock have to be if you want the portfolio risk to equal that of the overall market? A. 1.25 B. 1.50 C. 1.75 D. 2.25 E. 2.75
Answer #1

Weight of each=(1/3)

Portfolio beta=Respective beta*Respective investment weight

1=(1/3*0.75)+(1/3*0)+(1/3*Beta of other stock)[Beta of market=1;Beta of risk-free assets=0]

1=0.25+(1/3*Beta of other stock)

Beta of other stock=(1-0.25)*3

=2.25